# Simultaneous Uniformization and Algebraic Correspondences

### v1

@@ 3. Algebraic correspondences uniformizing two genus zero orbifolds
The passage from a genus zero orbifold group to its factor Bowen-Series map can be thought of as a ‘forgetful procedure’ from an invertible dynamical system to a non-invertible one. However, as explained in the previous section, conformal matings of factor Bowen-Series maps with polynomials can in fact be promoted to algebraic correspondences where the complete dynamical structure of the groups are recovered.
@@ 3.1. Conformal mating of factor Bowen-Series maps
The existence of the desired conformal mating is a consequence of [LMMN20, Theorem 5.2]. We include the mating construction for completeness and future reference.
Let $P_{0}(z)$ be the map $z\mapsto z^{d}$, where $d$ is the common degree of $A_{1},A_{2}$ on $\mathbb{S}^{1}$. By the proof of [MM23c, Lemma 3.4], each $A_{j}$, $j\in\{1,2\}$, admits a Markov partition satisfying conditions (4.1) and (4.2) of [LMMN20, Theorem 5.2]. Moreover, each periodic break-point of its piecewise analytic definition is symmetrically parabolic (cf. [LMMN20, Definition 4.6, Remark 4.7]). By [LMMN20, Theorem 4.13], the circle homeomorphisms $\mathfrak{g}_{j}$, that conjugate $P_{0}$ to $A_{j}$, $j\in\{1,2\}$, extend continuously to David homeomorphisms of $\mathbb{D}$.
@@
where $\mathbb{D}^{*}:=\widehat{\mathbb{C}}\setminus\overline{\mathbb{D}}$. The two definitions agree on $\mathbb{S}^{1}$. The domain of definition $\mathrm{Dom}(\widetilde{F})$ of the topological mating is ^$\widehat{\mathbb{C}}\setminus\left(\mathfrak{g}_{1}^{-1}(\operatorname{Int}{\mathcal{H}}_{1})\cup\widetilde{\mathfrak{g}_{2}}^{-1}(\operatorname{Int}{\mathcal{H}}_{2})\right)$ (see Figure 3.1).
@@
Next, we define an $\widetilde{F}$-invariant David (Beltrami) coefficient $\mu$ on $\widehat{\mathbb{C}}$ as follows (see [LMMN20, §2] for background on David homeomorphisms and David coefficients). In $\mathbb{D}$, we let $\mu$ be the pullback of the standard complex structure under the David homeomorphism $\mathfrak{g}_{1}$. In ^$\widehat{\mathbb{C}}\setminus\overline{\mathbb{D}}$, we let $\mu$ be the pullback of the standard complex structure under map $\widetilde{\mathfrak{g}}_{2}$. Then, $\mu$ is a David coefficient on $\widehat{\mathbb{C}}$. Since $A_{1},A_{2}$ are holomorphic, it follows that $\mu$ is $\widetilde{F}$-invariant.
@@
$F:=H\circ\widetilde{F}\circ H^{-1}:\mathrm{Dom}(F):=H\left(\mathrm{Dom}(\widetilde{F})\right)\to\widehat{\mathbb{C}}.$
By [LMMN20, Theorem 2.2], the maps $\mathfrak{X}_{1}:=H\circ\mathfrak{g}_{1}^{-1}:\mathbb{D}\to\widehat{\mathbb{C}}$ and $\mathfrak{X}_{2}:=H\circ\widetilde{\mathfrak{g}}_{2}^{-1}:\mathbb{D}\to\widehat{\mathbb{C}}$ are conformal. Hence, $F$ is holomorphic on $\operatorname{Int}{(\mathrm{Dom}(F))}\setminus\Lambda$, where $\Lambda:=H(\mathbb{S}^{1})$. Further, $F$ extends to a homeomorphism in a neighborhood of $\Lambda$, pinched at finitely many points, such that this extension is conformal outside $\Lambda$. Since David circles are locally conformally removable (cf. [LMMN20, Theorem 2.8]), it follows the above local extensions are conformal. Hence, $F$ is holomorphic on $\operatorname{Int}{\mathrm{Dom}(F)}$. It is readily checked that $\mathfrak{X}_{1},\mathfrak{X}_{2}$ are the desired mating conjugacies.
@@
[…]
@@ 3.2. Correspondence uniformizing a pair of genus zero orbifolds
For notational convenience, we denote the domain of $R_{j}$ by ^$\widehat{\mathbb{C}}_{j}$. Consider the disjoint union
@@ 4.1. Hecke orbifold and punctured spheres
Let $\Gamma_{1}\cong\mathbb{Z}/2\mathbb{Z}*\mathbb{Z}/2n\mathbb{Z}$ be the Fuchsian group such that $\mathbb{D}/\Gamma_{1}$ is the Hecke orbifold $\Sigma_{1}$; i.e., the genus zero orbifold with one puncture, an order $2$ orbifold point and an order $2n$ orbifold point, for $n\geq 2$. Let $A_{1}:\overline{\mathbb{D}}\setminus\operatorname{Int}{\mathcal{H}_{1}}\to\overline{\mathbb{D}}$ be the factor Bowen-Series map of $\Sigma_{1}$. The map $A_{1}$ restricts to a degree $12n-$ 1 covering of $\mathbb{S}^{1}$, and has a unique critical point, of multiplicity $12n-$ 1 (see Figure 4.1).
@@
Further, let $\Gamma_{2}\in\mathrm{Teich}(S_{0,n+1})$. We equip $\Gamma_{2}$ with the fundamental domain $\Pi_{2}$ described in Section 2.2, and set $\mathcal{H}_{2}:=\Pi_{2}$. Note that the factor Bowen-Series map of $\Gamma_{2}$ is the usual Bowen-Series map $A_{2}:\overline{\mathbb{D}}\setminus\operatorname{Int}{\mathcal{H}_{2}}\to\overline{\mathbb{D}}$. We denote the factor Bowen-Series map of any marked group $\Gamma\in\mathrm{Teich}(\Gamma_{2})\equiv\mathrm{Teich}(S_{0,n+1})$ by $A_{\Gamma}:\overline{\mathbb{D}}\setminus\operatorname{Int}{\mathcal{H}_{\Gamma}}\to\overline{\mathbb{D}}$. The map $A_{\Gamma}$ is a piecewise Möbius degree $12n-$ 1 circle covering (see Figure 4.2).
@@
We note that the Teichmüller space of $\Sigma_{2}$ has complex dimension $n-2$.
@@ 4.2. Conformal matings and associated correspondences
Figure 4.3. The shaded region is the domain of the topological mating between the factor Bowen-Series map $A_{1}$ of the $(2,6,\infty)$ genus zero orbifold and a Bowen-Series map $A_{\Gamma}$ of a four times punctured sphere group $\Gamma$. Here $\mathfrak{g}_{\Gamma}$ (respectively, $\widetilde{\mathfrak{g}}_{1}$) is a David homeomorphism from $\overline{\mathbb{D}}$ (respectively, from ^$\widehat{\mathbb{C}}\setminus\mathbb{D}$) onto $\overline{\mathbb{D}}$ that conjugates $z^{5}$ to $A_{\Gamma}$ (respectively, to $A_{1}$) on $\mathbb{S}^{1}$. The unique critical point and some of the fixed points and $2$-cycles of the mating are marked.
@@
    $\eta(\mathfrak{D})=\widehat{\mathbb{C}}\setminus\overline{\mathfrak{D}}$, ±$\pm 1\in\partial\mathfrak{D}$,
@@ 4.3. Explicit description of the correspondences
Note that $A_{1}$ has a unique critical point. This critical point has multiplicity $12n-$ 1, and has 00 as its associated critical value. Hence, the conformal mating $F$ also has a unique critical point $\boldsymbol{c}$, of multiplicity $12n-$ 1. We pre-compose $R$ with $N\in C(\eta)$ and post-compose $R$ with $M\in\mathrm{PSL}_{2}(\mathbb{C})$ such that $\boldsymbol{c}=0,F(0)=\infty,$ and $0\in\mathfrak{D}$ with $R(0)=0$. With these normalizations, Relation 3 implies that $R$ maps $\infty$ to $\infty$ with local degree $2n$, and hence $R$ is a degree $2n$ polynomial.
The ideal boundary points of $\mathcal{H}_{\Gamma}$ which are not fixed by $A_{\Gamma}$ form $(n-1)$ two-cycles. Hence, $F$ has $(n-1)$ two-cycles on $\partial\mathcal{D}$, such that $F$ does not extend analytically to neighborhoods of these points (see Figure 4.4). The above observation and Relation 3 imply that there exist $c_{1},\cdots,c_{n-1}\in\partial\mathfrak{D}$ such that for $j\in\{1,\cdots,n-1\}$, the points $c_{j},\eta(c_{j})$ are critical points of $R$, and they map under $R$ to these $(n-1)$ two-cycles of $F$.
@@
We also note that the unique cut-point $\boldsymbol{x}$ of $\partial\mathcal{D}$ is also a fixed point of $F$. Since $\partial\mathcal{D}$ is topologically the wedge of two circles, there exist $\beta,\beta^{\prime}\in\partial\mathfrak{D}$ ($\beta\neq\beta^{\prime}$) such that $R(\beta)=R(\beta^{\prime})=\boldsymbol{x}$ (see Figure 4.4). By Relation 3, we have $\beta^{\prime}=\eta(\beta)$. Further, the parabolic behavior of $A_{1},A_{\Gamma}$ at $1$ translates to the fact that the two branches of $F$ at $\boldsymbol{x}$ extend locally as tangent-to-identity parabolic germs. Hence, one has $R^{\prime}(\eta(\beta))\cdot\eta^{\prime}(\beta)=R^{\prime}(\beta)$. Therefore, the coefficients of $R$ satisfy the equation
$\mathrm{Res}_{z}(R(z)-R(\eta(z)),R^{\prime}(\eta(z))\cdot\eta^{\prime}(z)-R^{\prime}(z))=0,$
where $\mathrm{Res}_{z}(P_{1},P_{2})$ is the resultant of two univariate polynomials $P_{1},P_{2}\in\mathbb{C}[a_{2},\cdots,a_{n}][z]$. Therefore, the rational map $R$ (normalized as above) lies on the $(n-2)$-dimensional algebraic variety
(4.2) $\Big{\{}\left(a_{2},\cdots,a_{n}\right)\in\mathbb{C}^{n-1}:\mathrm{Res}_{z}\left(R(z)-R(\eta(z)),R^{\prime}(\eta(z))\cdot\eta^{\prime}(z)-R^{\prime}(z)\right)=0\Big{\}},$
@@ 4.4. Recovering marked groups from correspondences
Since $\infty\in\mathfrak{X}_{1}(\mathcal{H}_{1})$ and $R_{\Gamma}$ is a polynomial, it follows that $\mathcal{T}_{1}$ is a simply connected domain. In particular, $\mathcal{T}_{1}$ is the unique component of ^$\widehat{\mathbb{C}}\setminus\widetilde{\Lambda}$ containing a critical value of $R_{\Gamma}$. Further, the action of $\mathfrak{C}_{\Gamma}$ on $\mathcal{T}_{1}$ is generated by $\eta$ and the $12n-$ 1 non-trivial deck transformations of the branched covering $R_{\Gamma}:\mathcal{T}_{1}\to\mathfrak{X}_{1}(\mathbb{D})$ (which is fully branched over $\infty$ and is unbranched otherwise). By the proof of [MM23c, Proposition 5.13], these conformal automorphisms of $\mathcal{T}_{1}$ generate a group that acts properly discontinuously on $\mathcal{T}_{1}$, and the corresponding quotient is biholomorphic to the Hecke orbifold $\Sigma_{1}$.
@@
Finally, construct a Jordan curve $\mathfrak{J}$ by connecting the finite critical points of $R_{\Gamma}$ and $\beta,\beta^{-1}$ by hyperbolic geodesics in the simply connected domains $U_{1},\cdots,U_{2n},\mathcal{T}_{1}$. By construction, $\eta(\mathfrak{J})=\mathfrak{J}$. It now follows from the relation between the conformal mating and correspondence planes and the normalization of $R_{\Gamma}$ that the map $R_{\Gamma}$ is univalent on one of the complementary components of $\mathfrak{J}$ (in this case, it is the component of ^$\widehat{\mathbb{C}}\setminus\mathfrak{J}$ containing the origin), and this component coincides with the domain $\mathfrak{D}$ such that $\overline{R_{\Gamma}(\mathfrak{D})}$ is the domain of definition of the conformal mating $F$ of $A_{1}$ and $A_{\Gamma}$. Thus, $R_{\Gamma}$ completely determines the conformal mating of $A_{1}$ and $A_{\Gamma}$. In particular, the Bowen-Series map of $\Gamma$ can be recovered from $R_{\Gamma}$. Since the Bowen-Series map of $\Gamma$ encodes a preferred generating set of $\Gamma$, we can recover the group $\Gamma$ and its preferred generating set from $R_{\Gamma}$.
As a consequence of the preceding discussion, we have the following:
@@
$\mathrm{Teich}(S_{0,n+1})\to\mathbb{C}^{n-1}$
$\Gamma\mapsto R_{\Gamma}\hskip 11.38092pt$
is injective.
4.5. Holomorphic embedding of $\mathrm{Teich}(S_{0,n+1})$ into a space of correspondences

### v2

@@ 3. Algebraic correspondences uniformizing two genus zero orbifolds
The passage from a genus zero orbifold group to its factor Bowen-Series map can be thought of as a ‘forgetful procedure’ from an invertible dynamical system to a non-invertible one. However, as explained in the previous section, conformal matings of factor Bowen-Series maps with polynomials can in fact be promoted to algebraic correspondences where the complete dynamical structure of the groups is recovered.
@@ 3.1. Conformal mating of factor Bowen-Series maps
The existence of the desired conformal mating is a consequence of [LMMN25, Theorem 5.2]. We include the mating construction for completeness and future reference.
Let $P_{0}(z)$ be the map $z\mapsto z^{d}$, where $d$ is the common degree of $A_{1},A_{2}$ on $\mathbb{S}^{1}$. By the proof of [MM23c, Lemma 3.4], each $A_{j}$, $j\in\{1,2\}$, admits a Markov partition satisfying conditions (4.1) and (4.2) of [LMMN25, Theorem 5.2]. Moreover, each periodic break-point of its piecewise analytic definition is symmetrically parabolic (cf. [LMMN25, Definition 4.6, Remark 4.7]). By [LMMN25, Theorem 4.13], the circle homeomorphisms $\mathfrak{g}_{j}$, that conjugate $P_{0}$ to $A_{j}$, $j\in\{1,2\}$, extend continuously to David homeomorphisms of $\mathbb{D}$.
@@
[…]
@@
Next, we define an $\widetilde{F}-$invariant David (Beltrami) coefficient $\mu$ on $\widehat{\mathbb{C}}$ as follows (see [LMMN25, §2] for background on David homeomorphisms and David coefficients). In $\mathbb{D}$, we let $\mu$ be the pullback of the standard complex structure under the David homeomorphism $\mathfrak{g}_{1}$. In $\widehat{\mathbb{C}}\setminus\overline{\mathbb{D}}$, we let $\mu$ be the pullback of the standard complex structure under map $\widetilde{\mathfrak{g}}_{2}$. Then, $\mu$ is a David coefficient on $\widehat{\mathbb{C}}$. Since $A_{1},A_{2}$ are holomorphic, it follows that $\mu$ is $\widetilde{F}-$invariant.
@@
[…]
By [LMMN25, Theorem 2.2], the maps $\mathfrak{X}_{1}:=H\circ\mathfrak{g}_{1}^{-1}:\mathbb{D}\to\widehat{\mathbb{C}}$ and $\mathfrak{X}_{2}:=H\circ\widetilde{\mathfrak{g}}_{2}^{-1}:\mathbb{D}\to\widehat{\mathbb{C}}$ are conformal. Hence, $F$ is holomorphic on $\operatorname{Int}{(\mathrm{Dom}(F))}\setminus\Lambda$, where $\Lambda:=H(\mathbb{S}^{1})$. Further, $F$ extends to a homeomorphism in a neighborhood of $\Lambda$, pinched at finitely many points, such that this extension is conformal outside $\Lambda$. Since David circles are locally conformally removable (cf. [LMMN25, Theorem 2.8]), it follows that the above local extensions are conformal. Hence, $F$ is holomorphic on $\operatorname{Int}{\mathrm{Dom}(F)}$. It is readily checked that $\mathfrak{X}_{1},\mathfrak{X}_{2}$ are the desired mating conjugacies.
@@
$\{\mathfrak{X}_{1}(\mathfrak{g}_{1}(w))=\mathfrak{X}_{2}(\mathfrak{g}_{2}(\overline{w})):w\ \textrm{ is an}\ r\textrm{-th root of unity}\}.$
@@ 3.2. Correspondence uniformizing a pair of genus zero orbifolds
For notational convenience, we denote the domain of $R_{j}$ by $\widehat{\mathbb{C}}_{j}$. Consider the disjoint union
@@ 4.1. Hecke orbifold and punctured spheres
Let $\Gamma_{1}\cong\mathbb{Z}/2\mathbb{Z}*\mathbb{Z}/2n\mathbb{Z}$ be the Fuchsian group such that $\mathbb{D}/\Gamma_{1}$ is the Hecke orbifold $\Sigma_{1}$; i.e., the genus zero orbifold with one puncture, an order $2$ orbifold point and an order $2n$ orbifold point, for $n\geq 2$. Let $A_{1}:\overline{\mathbb{D}}\setminus\operatorname{Int}{\mathcal{H}_{1}}\to\overline{\mathbb{D}}$ be the factor Bowen-Series map of $\Sigma_{1}$. The map $A_{1}$ restricts to a degree 2n12n-1 covering of $\mathbb{S}^{1}$, and has a unique critical point, of multiplicity 2n12n-1 (see Figure 4.1).
@@
Further, let $\Gamma_{2}\in\mathrm{Teich}(S_{0,n+1})$. We equip $\Gamma_{2}$ with the fundamental domain $\Pi_{2}$ described in Section 2.2, and set $\mathcal{H}_{2}:=\Pi_{2}$. Note that the factor Bowen-Series map of $\Gamma_{2}$ is the usual Bowen-Series map $A_{2}:\overline{\mathbb{D}}\setminus\operatorname{Int}{\mathcal{H}_{2}}\to\overline{\mathbb{D}}$. We denote the factor Bowen-Series map of any marked group $\Gamma\in\mathrm{Teich}(\Gamma_{2})\equiv\mathrm{Teich}(S_{0,n+1})$ by $A_{\Gamma}:\overline{\mathbb{D}}\setminus\operatorname{Int}{\mathcal{H}_{\Gamma}}\to\overline{\mathbb{D}}$. The map $A_{\Gamma}$ is a piecewise Möbius degree 2n12n-1 circle covering (see Figure 4.2).
@@
We note that the Teichmüller space of $\Sigma_{2}$ has complex dimension n2n-2.
@@ 4.2. Conformal matings and associated correspondences
Figure 4.3. The shaded region is the domain of the topological mating between the factor Bowen-Series map $A_{1}$ of the $(2,6,\infty)$ genus zero orbifold and a Bowen-Series map $A_{\Gamma}$ of a four times punctured sphere group $\Gamma$. Here $\mathfrak{g}_{\Gamma}$ (respectively, $\widetilde{\mathfrak{g}}_{1}$) is a David homeomorphism from $\overline{\mathbb{D}}$ (respectively, from $\widehat{\mathbb{C}}\setminus\mathbb{D}$) onto $\overline{\mathbb{D}}$ that conjugates $z^{5}$ to $A_{\Gamma}$ (respectively, to $A_{1}$) on $\mathbb{S}^{1}$. The unique critical point and some of the fixed points and 22-cycles of the mating are marked.
@@
    $\eta(\mathfrak{D})=\widehat{\mathbb{C}}\setminus\overline{\mathfrak{D}}$, $\pm 1\in\partial\mathfrak{D}$,
@@ 4.3. Explicit description of the correspondences
Note that $A_{1}$ has a unique critical point. This critical point has multiplicity 2n12n-1, and has 0 as its associated critical value. Hence, the conformal mating $F$ also has a unique critical point $\boldsymbol{c}$, of multiplicity 2n12n-1. We pre-compose $R$ with $N\in C(\eta)$ and post-compose $R$ with $M\in\mathrm{PSL}_{2}(\mathbb{C})$ such that $\boldsymbol{c}=0,F(0)=\infty,$ and $0\in\mathfrak{D}$ with $R(0)=0$. With these normalizations, Relation 3 implies that $R$ maps $\infty$ to $\infty$ with local degree $2n$, and hence $R$ is a degree $2n$ polynomial.
The ideal boundary points of $\mathcal{H}_{\Gamma}$ which are not fixed by $A_{\Gamma}$ form (n1)(n-1) two-cycles. Hence, $F$ has (n1)(n-1) two-cycles on $\partial\mathcal{D}$, such that $F$ does not extend analytically to neighborhoods of these points (see Figure 4.4). The above observation and Relation 3 imply that there exist $c_{1},\cdots,c_{n-1}\in\partial\mathfrak{D}$ such that for $j\in\{1,\cdots,n-1\}$, the points $c_{j},\eta(c_{j})$ are critical points of $R$, and they map under $R$ to these (n1)(n-1) two-cycles of $F$.
@@
We also note that the unique cut-point $\boldsymbol{x}$ of $\partial\mathcal{D}$ is also a fixed point of $F$. Since $\partial\mathcal{D}$ is topologically the wedge of two circles, there exist $\beta,\beta^{\prime}\in\partial\mathfrak{D}$ ($\beta\neq\beta^{\prime}$) such that $R(\beta)=R(\beta^{\prime})=\boldsymbol{x}$ (see Figure 4.4). By Relation 3, we have $\beta^{\prime}=\eta(\beta)$. Further, the parabolic behavior of $A_{1},A_{\Gamma}$ at $1$ translates to the fact that the two branches of $F$ at $\boldsymbol{x}$ extend locally as tangent-to-identity parabolic germs. Hence, the complex number $\beta$ satisfies the equations
$R(z)=R(\eta(z)),\quad\mathrm{and}\quad R^{\prime}(\eta(z))\cdot\eta^{\prime}(z)=R^{\prime}(z).$
Note that the above equations are also satisfied by $z=1$. Thus, we have that $\deg_{z}\left(\gcd\left(R(z)-R(\eta(z)),R^{\prime}(\eta(z))\cdot\eta^{\prime}(z)-R^{\prime}(z)\right)\right)\geq 2$. Hence the coefficients of $R$ satisfy the equation
$\mathrm{sRes}_{1}(R(z)-R(\eta(z)),R^{\prime}(\eta(z))\cdot\eta^{\prime}(z)-R^{\prime}(z))=0,$
where $\mathrm{sRes}_{j}(P_{1},P_{2})$ denotes the $j$-th principal subresultant coefficient of two univariate polynomials $P_{1},P_{2}\in\mathbb{C}[a_{2},\cdots,a_{n}][z]$ (cf. [BPC06, §4.2.2]).
Therefore, the rational map $R$ lies on the (n2)(n-2)-dimensional algebraic variety
(4.2) $\mathscr{V}:=\Big\{$ $\left(a_{2},\cdots,a_{n}\right)\in\mathbb{C}^{n-1}:$
$\mathrm{sRes}_{1}\left(R(z)-R(\eta(z)),R^{\prime}(\eta(z))\cdot\eta^{\prime}(z)-R^{\prime}(z)\right)=0\Big\},$
@@ 4.4. Recovering marked groups from correspondences
[…]
Since $\infty\in\mathfrak{X}_{1}(\mathcal{H}_{1})$ and $R_{\Gamma}$ is a polynomial, it follows that $\mathcal{T}_{1}$ is a simply connected domain. In particular, $\mathcal{T}_{1}$ is the unique component of $\widehat{\mathbb{C}}\setminus\widetilde{\Lambda}$ containing a critical value of $R_{\Gamma}$. Further, the action of $\mathfrak{C}_{\Gamma}$ on $\mathcal{T}_{1}$ is generated by $\eta$ and the 2n12n-1 non-trivial deck transformations of the branched covering $R_{\Gamma}:\mathcal{T}_{1}\to\mathfrak{X}_{1}(\mathbb{D})$ (which is fully branched over $\infty$ and is unbranched otherwise). By the proof of [MM23c, Proposition 5.13], these conformal automorphisms of $\mathcal{T}_{1}$ generate a group that acts properly discontinuously on $\mathcal{T}_{1}$, and the corresponding quotient is biholomorphic to the Hecke orbifold $\Sigma_{1}$.
@@
Finally, construct a Jordan curve $\mathfrak{J}$ by connecting the finite critical points of $R_{\Gamma}$ and $\beta,\beta^{-1}$ by hyperbolic geodesics in the simply connected domains $U_{1},\cdots,U_{2n},\mathcal{T}_{1}$. By construction, $\eta(\mathfrak{J})=\mathfrak{J}$. It now follows from the relation between the conformal mating and correspondence planes and the normalization of $R_{\Gamma}$ that the map $R_{\Gamma}$ is univalent on one of the complementary components of $\mathfrak{J}$ (in this case, it is the component of $\widehat{\mathbb{C}}\setminus\mathfrak{J}$ containing the origin), and this component coincides with the domain $\mathfrak{D}$ such that $\overline{R_{\Gamma}(\mathfrak{D})}$ is the domain of definition of the conformal mating $F$ of $A_{1}$ and $A_{\Gamma}$. Thus, $R_{\Gamma}$ completely determines the conformal mating of $A_{1}$ and $A_{\Gamma}$. In particular, the Bowen-Series map of $\Gamma$ can be recovered from $R_{\Gamma}$. Since the Bowen-Series map of $\Gamma$ encodes a preferred generating set of $\Gamma$, we can recover the group $\Gamma$ and its preferred generating set from $R_{\Gamma}$.
As a consequence of the preceding discussion, we have the following result.
@@
$\boldsymbol{\Psi}:\$$\mathrm{Teich}(S_{0,n+1})\to\mathscr{V}\subset\mathbb{C}^{n-1}$
$\hskip 56.9055pt\Gamma\mapsto R_{\Gamma}$
is injective. Here, $\mathscr{V}$ is the (n2)(n-2)-dimensional algebraic variety defined by Equation (4.2).
Figure 4.4: The conformal mating $F$ is defined on the ‘pinched annulus’ $\widehat{\mathbb{C}}\setminus\operatorname{Int}{\left(\mathfrak{X}_{\Gamma}(\mathcal{H}_{\Gamma})\cup\mathfrak{X}_{1}(\mathcal{H}_{1})\right)}$, where $\mathfrak{X}_{\Gamma}(\mathcal{H}_{\Gamma})$ is the bounded region enclosed by the red hexagon, and $\mathfrak{X}_{1}(\mathcal{H}_{1})$ is the region outside the blue monogon. The black fractal Jordan curve is the limit set $\Lambda$ of the conformal mating $F$. The bounded complementary component of $\Lambda$ is $\mathfrak{X}_{\Gamma}(\mathbb{D})$, while the unbounded complementary component is $\mathfrak{X}_{1}(\mathbb{D})$. The first preimages of $\mathfrak{X}_{\Gamma}(\mathcal{H}_{\Gamma})$, $\mathfrak{X}_{1}(\mathcal{H}_{1})$ under $F$ are also shown.
The set $\mathfrak{X}_{1}(\mathcal{H}_{1})$ (respectively, $\mathfrak{X}_{1}(\mathbb{D})$) has a connected preimage under $R_{\Gamma}$ in the correspondence plane, and it maps onto $\mathfrak{X}_{1}(\mathcal{H}_{1})$ (respectively, $\mathfrak{X}_{1}(\mathbb{D})$) as a $6:1$ branched cover ramified only at $\infty$. On the other hand, $R_{\Gamma}^{-1}(\mathfrak{X}_{\Gamma}(\mathcal{H}_{\Gamma}))$ (respectively, $R_{\Gamma}^{-1}(\mathfrak{X}_{\Gamma}(\mathbb{D}))$) has six connected components, each of which maps conformally onto $\mathfrak{X}_{\Gamma}(\mathcal{H}_{\Gamma})$ (respectively, $\mathfrak{X}_{\Gamma}(\mathbb{D})$). The limit set $\widetilde{\Lambda}=R_{\Gamma}^{-1}(\Lambda)$ of the correspondence $\mathfrak{C}_{\Gamma}$ is a chain of six Jordan curves that divides the sphere into the fully invariant sets $\mathcal{T}_{\Gamma}=R_{\Gamma}^{-1}(\mathfrak{X}_{\Gamma}(\mathbb{D}))\quad\mathrm{and}$ and $\mathcal{T}_{1}=R_{\Gamma}^{-1}(\mathfrak{X}_{1}(\mathbb{D}))$.
4.5. Holomorphic injection of $\mathrm{Teich}(S_{0,n+1})$ into a space of correspondences